OpenStudy (anonymous):

find a system of inequalities to describe the given region (2,2)(-2,2)(-2,-2)(2,-2)

OpenStudy (anonymous):

\[-2\le x \le 2, -2\le y \le2\]

OpenStudy (anonymous):

i think this is right.... not 100% sure though.